OpenStudy (anonymous):

If the pressure is constant and the temperature increased from 20°C to 25°C, what was the initial volume if the final volume of the temperature increase is 45.5 L? Solve and state what formular and law (Charles or Boyle) is used Thanks in advance! \(\huge\ddot\smile\)

OpenStudy (fifciol):

Charles law if pressure is constant: \[\frac{ pV_1 }{ T_1 }=\frac{ pV_2 }{T_2 }\rightarrow V_1=\frac{ T_1 }{ T_2 }V_2\] Remember to convert celsius to kelwin 1Kelwin= T in celius +273

OpenStudy (anonymous):

Thank you!!